OpenStudy (technodynamic):

Suppose f(x)=sqrt32+4x^2 Calculate the exact value(s) of x when f(x)=6.If there is more than one such value, enter all possible values as a comma separated list.

OpenStudy (anonymous):

ah! its different than the way you wrote it before .. 6 = sqrt(32) + 4x^2 sqrt(32) = sqrt(16*2) = 4*sqrt(2) 1.5 -sqrt(2) = x^2 x = +/- sqrt(1.5-sqrt(2))